News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 9 August 2002 2.15pm EDT

During the last week, Chandra completed the observing schedule as planned.

Observations included a target of opportunity, GX 1+4, coordinated with RXTE; an observation of Uranus; a simultaneous ACIS/HRC observation in which HRC-I observed RXJ1856.5-3754 while ACIS observed the background; and two observations coordinated with ground telescopes.

Chandra passed through eclipses 1 and 2 of the 2002 Fall season on August 7 and 9 with nominal power and thermal performance.

A routine real-time procedure was completed on August 7 to update the ephemeris.

A press release was issued on August 8 featuring a composite X-ray, radio and optical image of Centaurus A (http://chandra.harvard.edu/press/02_releases/press_080702.html) that is believed to provide evidence of a giant eruption in the center of the galaxy about 10 million years ago.

The observing schedule for next week is shown below.

------------------------------------------
M31                     ACIS-I       Aug 11
TEV J2032+4130          ACIS-I
GROTH-WESTPHAL FIELD    ACIS-I
Radiation Belts                      Aug 12
GROTH-WESTPHAL FIELD    ACIS-I
Radiation Belts                      Aug 14
IM PEG                  ACIS-S/HETG  Aug 15
PSS 2311+1644           ACIS-S
BETA CETI               HRC-S/LETG
4U 1850-087             ACIS-S/LETG  Aug 16
1E 1740.7-2942          ACIS-S       Aug 17
Radiation Belts
2S 1845-024             ACIS-S       Aug 18
AQL X-1                 ACIS-S
A2065                   ACIS-I
NGC7052                 ACIS-S
A2409                   ACIS-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports